OpenStudy (anonymous):

Part 1: Explain, in complete sentences, how you would use the substitution method to solve the following system of equations. Part 2: Provide the solution to the system. 5x – y = –2 10x – 3y = –7

OpenStudy (anonymous):

First, solve for y in each equation. 5x - y = -2 -y = -2 - 5x y = 5x + 2 10x-3y = -7 10x = -7 + 3y 10x+7 = 3y y = (1/3)(10x+7) Next, because each equation is set to equal y, they must be equal to each other. For example, if y = w and y = x, w must be equal to x. Solve for x. y = y 5x + 2 = (1/3)(10x + 7) 15x+6 = 10x+7 5x = 1 x=1/5 Now that you have a value of x, you can substitute x in for any of the two equations for the y-value. 5x - y = -2 5(1/5) - y = -2 1 + 2 = y y = 3 The solution to this system of equations is: (1/5, 3)

OpenStudy (anonymous):

thank you check out my other problems maybe you could help me PLEASE!!!